Rewrite the equation 2x – 3y = –6 as a function of x
Answer by mala_kumarkos@yahoo.com(10)


2x-3y=-6
2x+6=3y
3y=2x+6
y=2x+6/3
therefore y = f(x)=(2x/3+6/3)